A steel rod on a bridge must be made to withstand a pull of 5000 lbs. Find the diameter of


the rod assuming the factor safety of 5 and the ultimate stress of 64000 psi. (15 pts)

1
Expert's answer
2021-12-21T12:16:52-0500

With the safety factor of 5, the rod must withstand a load of 5·5000=25000 lbs. The cross-sectional area of such a rod should be


"A=F/\sigma=2.52·10^{-4}\text{ m}^2,\\\space\\
d=2\sqrt{\frac A\pi}=0.018 \text{ m},"

or 1.8 cm, or 0.709 in.